OpenStudy (anonymous):

I have several issue's I'm having with algebra, can anyone help me with simplifying some expressions? Simplify the expression: (-4)(-x) I think it's -4X but im not sure. my parents have tried to help me with this but we're still having a rough time. is anyone willing to help?

OpenStudy (texaschic101):

-4 * -x = 4x negative * negative = positive

OpenStudy (anonymous):

negative times negative is positive, and 4 times x is 4x, so what do you think?

OpenStudy (anonymous):

ah... ok. my mom helped me do this lol this is sad. alright, can you guys help me with a few more?

OpenStudy (texaschic101):

lets see what ya got

OpenStudy (anonymous):

Simplify the expression: (-2)(-y)(y) we came up with -2Y x Y

OpenStudy (anonymous):

what's y*y?

OpenStudy (anonymous):

but now you're elling me that would be a positive. but are the placement of the letters even correct? I don't even know

OpenStudy (anonymous):

uhm... y? XD

OpenStudy (texaschic101):

(-2)(-y)(y) = 2y * y = 2y^2

OpenStudy (anonymous):

y*y is y^2

OpenStudy (anonymous):

so 2y^2? thank you

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

yay! ok next one is Simplify the expression: 7(x-5) we came up with 7X-5

OpenStudy (anonymous):

you know what the distributive property is?

OpenStudy (texaschic101):

7(x - 5) --- distribute through the parenthesis all that means is multiply everything in the parenthesis by the number outside of the parenthesis, which is 7. 7(x - 5) = 7 * x - 7 * 5 = 7x - 35

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

The distributive property lets you multiply the 7 across whatever is in the parentheses you would apply the distributive property to this: (7*x) + (7*-5) so... 7x + -35

OpenStudy (anonymous):

which is 7x - 35

OpenStudy (texaschic101):

you multiplied the 7 times x, but you did not multiply the 7 times 5. You have to multiply the 7 by EVERYTHING in the parenthesis 7(x - 5) = 7 * x = 7x 7 * -5 = -35 all together = 7x - 35

OpenStudy (anonymous):

Thank you both so much! alright, next one Simplify the expression: (3-x)(4) we came up with -3x*4 but now you're telling me... so.. -3 times 4.. so...

OpenStudy (anonymous):

that's so off.

OpenStudy (anonymous):

alright... 12... 12-x?

OpenStudy (texaschic101):

that can be written as 4(3 - x) multiply 4 by EVERYTHING in parenthesis 4 * 3 = 12 4 * -x = -4x put them together : 12 - 4x or -4x + 12

OpenStudy (anonymous):

4 times x too..

OpenStudy (anonymous):

12 -x^4? XD

OpenStudy (anonymous):

lol ok so... it's -4x plus 12... so -16x?

OpenStudy (anonymous):

I'm sorry. I'm terrible with math and just the principles of all this too.

OpenStudy (texaschic101):

no... 4(3 - x) = 4 * 3 = 12 4 * -x = -4x together = 12 - 4x or -4x + 12

OpenStudy (anonymous):

hmmm... the 4 is multiplied by the x not put as a power... it would actually be -4x... so you would have -4x + 12... and you can't simplify any further

OpenStudy (anonymous):

its ok :D

OpenStudy (anonymous):

>.< thanks for being willing to help. I'm so stuck with all this. here's the next one.. Ima try to figure these out too lol Simplify the expression: 3(2-4x)

OpenStudy (texaschic101):

multiply 3 by EVERYTHING in parenthesis

OpenStudy (anonymous):

so 3*2, and 3*-4x ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

yep

OpenStudy (anonymous):

so... simplifying farther...

OpenStudy (texaschic101):

what did you get when you multiplied ?

OpenStudy (anonymous):

3*2 is 6 and 3*-4x is -12x...?

OpenStudy (texaschic101):

which equals 6 - 12x or -12x + 6

OpenStudy (anonymous):

which one should i put in?

OpenStudy (anonymous):

both are correct...

OpenStudy (texaschic101):

I would put 6 - 12x....just put them in the order you multiplied

OpenStudy (anonymous):

alright, i am. thank you. but my school is touchy so just making sure >.< next one is Simplify the expression: -(10-7x)

OpenStudy (texaschic101):

but it does not really matter because they are both correct

OpenStudy (anonymous):

have you tried khan academy? there are lots of helpful videos...

OpenStudy (anonymous):

so... 10-7 is 3, so -3x?

OpenStudy (anonymous):

actually... you can't subtract it...

OpenStudy (anonymous):

i have looked at kahn academy, but when i found this it's nice to have live help that explains it to me.... so thank you guys for helping me out with this. it's easier to have someone explain to me. I cant wait to be in my new online school where there will be a teacher and a tutor.

OpenStudy (texaschic101):

ok...when you have a sign without the number outside of the parenthesis, you can put a 1, if it makes it easier. -(10 - 7x) can also be written as -1(10 - 7x) now just multiply everything in the parenthesis by -1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-1*10 is -10, and -1*-7x is -7x right? so it would be... -10 - (-7x) ?

OpenStudy (texaschic101):

not quite...-1 * 10 = -10 but -1 * -7x = 7x (negative * negative = positive) giving us : -10 + 7x

OpenStudy (texaschic101):

it would be easier for you to learn these rules... this only applies to multiplication/division ... if the signs are alike (either both neg or pos), then the ending sign will be positive if the signs are different, the ending result will be negative DO NOT GET THESE RULES CONFUSED WITH ADDING/SUBTRACTING....there are different rules for those

OpenStudy (anonymous):

ohhhh i see! well i was close..

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

here's the enxt one.. Simplify the expression: 40x ÷ (1/8)

OpenStudy (anonymous):

division....

OpenStudy (anonymous):

ok... so what did you get?

OpenStudy (anonymous):

whenever you divide by a fraction, you multiply by the reciprocal. do you know what the reciprocal is?

OpenStudy (anonymous):

So, dividing by 1/8 becomes multiply by 8/1. Which is 8. So here: 40x * 8 320x

OpenStudy (anonymous):

reciprocal is the flip of the fraction: 1/8 becomes 8

OpenStudy (anonymous):

@darkest_midnight you get it?

OpenStudy (anonymous):

not really... 40x divided by the reciprocal of one which is 8, (i do understand reciprocals)

OpenStudy (anonymous):

so 40 divided by 8 is 5,

OpenStudy (anonymous):

1/8 reciprocal is 8 so after you reciprocal it... you multiply instead of divide...

OpenStudy (texaschic101):

true...when dividing fractions, " flip " what you are dividing by, then just multiply Example : 2/3 divided by 1/4 = 2/3 * 4/1 8 divided by 1/5 = 8 * 5/1 3/7 divided by 2 = 3/7 * 1/2 just remember to " flip " what you are dividing by, not the other number

OpenStudy (anonymous):

so 40 times 8 is 320

OpenStudy (anonymous):

ok so multiplication is 320, now what do we do with that number?

OpenStudy (anonymous):

its 320 times x, right? so what would that be?

OpenStudy (anonymous):

so 320x

OpenStudy (anonymous):

wow. thanks...

OpenStudy (texaschic101):

there's you answer darkest_midnight....320x

OpenStudy (anonymous):

you're welcome :D

OpenStudy (anonymous):

this one's really fun... Simplify the expression: (c/3) ÷ (5/6)

OpenStudy (anonymous):

ooooh... ok fun!

OpenStudy (texaschic101):

remember to " flip " what you are dividing by, then just multiply

OpenStudy (anonymous):

so which do you think you reciprocate?

OpenStudy (anonymous):

so 6*3 and c*5?

OpenStudy (anonymous):

only flip the one you are dividing by...

OpenStudy (anonymous):

(c/3) * (6/5)

OpenStudy (anonymous):

you are dividing by 5/6 so reciprocate that then multiply

OpenStudy (anonymous):

so (6*c)/(5*3)

OpenStudy (anonymous):

which is....?

OpenStudy (anonymous):

well that would be 6c/15?

OpenStudy (anonymous):

yeah! and that simplifies to?

OpenStudy (anonymous):

6/15 simplifies to?

OpenStudy (anonymous):

1/4 x

OpenStudy (anonymous):

1/4x?

OpenStudy (anonymous):

you understand?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

not completely... im trying though ^-^

OpenStudy (anonymous):

here... lets see if i can get this next one... Simplify the expression: (8x+6)/2

OpenStudy (anonymous):

yeah! you're one of very few askers who put effort into their work here :D

OpenStudy (anonymous):

so first... this is distributive division...

OpenStudy (anonymous):

which means that you divide EVERYTHING by 2... so

OpenStudy (anonymous):

8x divided by 2 is 4...

OpenStudy (anonymous):

6/2 is 3....